# Agentic Context Engineering: Evolving Contexts for Self-Improving Language Models

### v1

@@ 1 Introduction
Modern AI applications based on large language models (LLMs), such as LLM agents [52, 49] and compound AI systems [55], increasingly depend on context adaptation. Instead of modifying model weights, context adaptation improves performance after model training by incorporating clarified instructions, structured reasoning steps, or domain-specific input formats directly into the model’s inputs. Contexts underpin many AI system components, including system prompts that guide downstream tasks [36, 4], memory that carries past facts and experiences [41, 48], and factual evidence that reduces hallucination and supplements knowledge [6].
@@
Adapting through contexts rather than weights offers several key advantages. Contexts are interpretable and explainable for users and developers [47, 45], allow rapid integration of new knowledge at runtime [27, 7], and can be shared across models or modules in a compound system [23]. Meanwhile, advances in long-context LLMs [39] and context-efficient inference such as KV cache reuse [17, 51] are making context-based approaches increasingly practical for deployment. As a result, context adaptation is emerging as a central paradigm for building capable, scalable, and self-improving AI systems.
@@
Despite this progress, existing approaches to context adaptation face two key limitations. First, a brevity bias: many prompt optimizers prioritize concise, broadly applicable instructions over comprehensive accumulation. For example, GEPA [4] highlights brevity as a strength, but such abstraction can omit domain-specific heuristics, tool-use guidelines, or common failure modes that matter in practice [16]. This objective aligns with validation metrics in some settings, but often fails to capture the detailed strategies required by agents and knowledge-intensive applications. Second, context collapse: methods that rely on monolithic rewriting by an LLM often degrade into shorter, less informative summaries over time, causing sharp performance declines (Figure 2). In domains such as interactive agents [43, 38, 57], domain-specific programming [53, 56], and financial or legal analysis [33, 18, 44], strong performance depends on retaining detailed, task-specific knowledge rather than compressing it away.
@@
As applications such as agents and knowledge-intensive reasoning demand greater reliability, recent work has shifted toward saturating contexts with abundant, potentially useful information [22, 12, 11], enabled by advances in long-context LLMs [39, 34]. We argue that contexts should function not as concise summaries, but as comprehensive, evolving playbooks—detailed, inclusive, and rich with domain insights. Unlike humans, who often benefit from concise generalization, LLMs are more effective when provided with long, detailed contexts and can distill relevance autonomously [22, 31, 41]. Thus, instead of compressing away domain-specific heuristics and tactics, contexts should preserve them, allowing the model to decide what matters at inference time.
@@
To address these limitations, we introduce ACE (Agentic Context Engineering), a framework for comprehensive context adaptation in both offline settings (e.g., system prompt optimization) and online settings (e.g., test-time memory adaptation). Rather than compressing contexts into distilled summaries, ACE treats them as evolving playbooks that accumulate and organize strategies over time. Building on the agentic architecture of Dynamic Cheatsheet [41], ACE incorporates a modular workflow of generation, reflection, and curation, while adding structured, incremental updates guided by a grow-and-refine principle. This design preserves detailed, domain-specific knowledge, prevents context collapse, and yields contexts that remain comprehensive and scalable throughout adaptation.
@@
We evaluate ACE on two categories of LLM applications that most benefit from comprehensive, evolving contexts: (1) agents [43], which require multi-turn reasoning, tool use, and environment interaction, where accumulated strategies can be reused across episodes; and (2) domain-specific benchmarks, which demand specialized tactics and knowledge, where we focus on financial analysis [33, 44]. Our key findings are:
@@
    On the AppWorld benchmark leaderboard [5], ACE matches the top-ranked production-level agent IBM-CUGA [35] (powered by GPT-4.1) on average and surpasses it on the harder test-challenge split, while using a smaller open-source model (DeepSeek-V3.1).
@@ 2.1 Context Adaptation
Context adaptation (or context engineering) refers to methods that improve model behavior by constructing or modifying inputs to an LLM, rather than altering its weights. The current state of the art leverages natural language feedback [40, 54, 4]. In this paradigm, a language model inspects the current context along with signals such as execution traces, reasoning steps, or validation results, and generates natural language feedback on how the context should be revised. This feedback is then incorporated into the context, enabling iterative adaptation. Representative methods include Reflexion [40], which reflects on failures to improve agent planning; TextGrad [54], which optimizes prompts via gradient-like textual feedback; GEPA [4], which refines prompts iteratively based on execution traces and achieves strong performance, even surpassing reinforcement learning approaches in some settings; and Dynamic Cheatsheet [41], which constructs an external memory that accumulates strategies and lessons from past successes and failures during inference. These natural language feedback methods represent a major advance, offering flexible and interpretable signals for improving LLM systems beyond weight updates.
@@ 2.2 Limitations of Existing Context Adaptation Methods
A recurring limitation of context adaptation methods is brevity bias: the tendency of optimization to collapse toward short, generic prompts. Gao et al. [16] document this effect in prompt optimization for test generation, where iterative methods repeatedly produced near-identical instructions (e.g., "Create unit tests to ensure methods behave as expected"), sacrificing diversity and omitting domain-specific detail. This convergence not only narrows the search space but also propagates recurring errors across iterations, since optimized prompts often inherit the same faults as their seeds. More broadly, such bias undermines performance in domains that demand detailed, context-rich guidance—such as multi-step agents, program synthesis, or knowledge-intensive reasoning—where success hinges on accumulating rather than compressing task-specific insights.
@@
In a case study on the AppWorld benchmark [43], we observe a phenomenon we call context collapse, which arises when an LLM is tasked with fully rewriting the accumulated context at each adaptation step. As the context grows large, the model tends to compress it into much shorter, less informative summaries, causing a dramatic loss of information. For instance, at step 60 the context contained 18,282 tokens and achieved an accuracy of 66.7, but at the very next step it collapsed to just 122 tokens, with accuracy dropping to 57.1—worse than the baseline accuracy of 63.7 without adaptation. While we highlight this through Dynamic Cheatsheet [41], the issue is not specific to that method; rather, it reflects a fundamental risk of end-to-end context rewriting with LLMs, where accumulated knowledge can be abruptly erased instead of preserved.
@@ 3 Agentic Context Engineering (ACE)
We present ACE (Agentic Context Engineering), a framework for scalable and efficient context adaptation in both offline (e.g., system prompt optimization) and online (e.g., test-time memory adaptation) scenarios. Instead of condensing knowledge into terse summaries or static instructions, ACE treats contexts as evolving playbooks that continuously accumulate, refine, and organize strategies over time. Building on the agentic design of Dynamic Cheatsheet [41], ACE introduces a structured division of labor across three roles (Figure 4): the Generator, which produces reasoning trajectories; the Reflector, which distills concrete insights from successes and errors; and the Curator, which integrates these insights into structured context updates. This mirrors how humans learn—experimenting, reflecting, and consolidating—while avoiding the bottleneck of overloading a single model with all responsibilities.
@@ 3.1 Incremental Delta Updates
A core design principle of ACE is to represent context as a collection of structured, itemized bullets, rather than a single monolithic prompt. The concept of a bullet is similar to the concept of a memory entry in LLM memory frameworks like Dynamic Cheatsheet [41] and A-MEM [48], but builds on top of that and consists of (1) metadata, including a unique identifier and counters tracking how often it was marked helpful or harmful; and (2) content, capturing a small unit such as a reusable strategy, domain concept, or common failure mode. When solving new problems, the Generator highlights which bullets were useful or misleading, providing feedback that guides the Reflector in proposing corrective updates.
@@ 4.1 Tasks and Datasets
    LLM Agent: AppWorld [43] is a suite of autonomous agent tasks involving API understanding, code generation, and environment interaction. It provides a realistic execution environment with common applications and APIs (e.g., email, file system) and tasks of two difficulty levels (normal and challenge). A public leaderboard [5] tracks performance, where, at the time of submission, the best system achieved only 60.3% average accuracy, highlighting the benchmark’s difficulty and realism.
@@
    Financial Analysis: FiNER [33] and Formula [44] test LLMs on financial reasoning tasks that rely on the eXtensible Business Reporting Language (XBRL). FiNER requires labeling tokens in XBRL financial documents with one of 139 fine-grained entity types, a key step for financial information extraction in regulated domains. Formula focuses on extracting values from structured XBRL filings and performing computations to answer financial queries, i.e., numerical reasoning.
@@ 4.2 Baselines and Methods
The base model is evaluated directly on each benchmark without any context engineering, using the default prompts provided by dataset authors. For AppWorld, we follow the official ReAct [52] implementation released by the benchmark authors, and build all other baselines and methods on top of this framework.
In-Context Learning (ICL) [3].
@@
MIPROv2 [36].
MIPROv2 is a popular prompt optimizer for LLM applications that works by jointly optimizing system instructions and in-context demonstrations via bayesian optimization. We use the official DSPy implementation [15], setting auto="heavy" to maximize optimization performance.
GEPA [4].
GEPA (Genetic-Pareto) is a sample-efficient prompt optimizer based on reflective prompt evolution. It collects execution traces (reasoning, tool calls, intermediate outputs) and applies natural-language reflection to diagnose errors, assign credit, and propose prompt updates. A genetic Pareto search maintains a frontier of high-performing prompts, mitigating local optima. Empirically, GEPA outperforms reinforcement learning methods such as GRPO and prompt optimizers like MIPROv2, achieving up to 10–20% higher accuracy with as much as 35× fewer rollouts. We use the official DSPy implementation [14], setting auto="heavy" to maximize optimization performance.
Dynamic Cheatsheet (DC) [41].
DC is a test-time learning approach that introduces an adaptive external memory of reusable strategies and code snippets. By continuously updating this memory with newly encountered inputs and outputs, DC enables models to accumulate knowledge and reuse it across tasks, often leading to substantial improvements over static prompting methods. A key advantage of DC is that it does not require ground-truth labels: the model can curate its own memory from its generations, making the method highly flexible and broadly applicable. We use the official implementation released by the authors [42] and set it to use the cumulative mode (DC-CU).
@@
ACE optimizes LLM contexts for both offline and online adaptation through an agentic context engineering framework. To ensure fairness, we use the same LLM for the Generator, Reflector, and Curator (non-thinking mode of DeepSeek-V3.1 [13]), preventing knowledge transfer from a stronger Reflector or Curator to a weaker Generator. This isolates the benefit of context construction itself. We adopt a batch size of 1 (constructing a delta context from each sample). We set the maximum number of Reflector refinement rounds and the maximum number of epochs in offline adaptation to 5.
@@ 4.3 Results on Agent Benchmark
Notably, on the latest AppWorld leaderboard (as of September 20, 2025; Figure 5), on average, ReAct + ACE (59.4%) matches the top-ranked IBM CUGA (60.3%), a production-level GPT-4.1–based agent [35], despite using the smaller open-source model DeepSeek-V3.1. With online adaptation, ReAct + ACE even surpasses IBM CUGA by 8.4% in TGC and 0.7% in SGC on the harder test-challenge split, underscoring the effectiveness of ACE in building comprehensive and self-evolving contexts for agents.
@@ 5 Discussion
Although ACE produces longer contexts than methods such as GEPA, this does not translate to linearly higher inference cost or GPU memory usage. Modern serving infrastructures are increasingly optimized for long-context workloads through techniques such as the reuse [17, 51], compression [32, 30], and offload [25] of KV cache. These mechanisms allow frequently reused context segments to be cached locally or remotely, avoiding repetitive and expensive prefill operations. Ongoing advances in ML systems suggest that the amortized cost of handling long contexts will continue to decrease, making context-rich approaches like ACE increasingly practical in deployment.
@@
Online and continuous learning are key research directions in machine learning for addressing issues like distribution shifts [24, 19] and limited training data [37, 21, 60]. ACE offers a flexible and efficient alternative to conventional model fine-tuning, as adapting contexts is generally cheaper than updating model weights [9, 26, 28, 20]. Moreover, because contexts are human-interpretable, ACE enables selective unlearning [10, 8, 29]—whether due to privacy or legal constraints [1, 2], or when outdated or incorrect information is identified by domain experts. These are promising directions for future work, where ACE could play a central role in advancing continuous and responsible learning.

### v2

@@ 1 Introduction
Modern AI applications based on large language models (LLMs), such as LLM agents [yao2023react, yang2024swe] and compound AI systems [zaharia2024compoundGS], increasingly depend on context adaptation. Instead of modifying model weights, context adaptation improves performance after model training by incorporating clarified instructions, structured reasoning steps, or domain-specific input formats directly into the model’s inputs. Contexts underpin many AI system components, including system prompts that guide downstream tasks [opsahl2024optimizing, agrawal2025gepa], memory that carries past facts and experiences [suzgun2025dynamic, xu2025mem], and factual evidence that reduces hallucination and supplements knowledge [asai2024self].
@@
Adapting through contexts rather than weights offers several key advantages. Contexts are interpretable and explainable for users and developers [wei2022chain, wang2022self], allow rapid integration of new knowledge at runtime [lewis2020retrieval, borgeaud2022improving], and can be shared across models or modules in a compound system [khot2022decomposed]. Meanwhile, advances in long-context LLMs [peng2023yarn] and context-efficient inference such as KV cache reuse [gim2024prompt, yao2025cacheblend] are making context-based approaches increasingly practical for deployment. As a result, context adaptation is emerging as a central paradigm for building capable, scalable, and self-improving AI systems.
@@
Despite this progress, existing approaches to context adaptation face two key limitations. First, a brevity bias: many prompt optimizers prioritize concise, broadly applicable instructions over comprehensive accumulation. For example, GEPA [agrawal2025gepa] highlights brevity as a strength, but such abstraction can omit domain-specific heuristics, tool-use guidelines, or common failure modes that matter in practice [gao2025prompt]. This objective aligns with validation metrics in some settings, but often fails to capture the detailed strategies required by agents and knowledge-intensive applications. Second, context collapse: methods that rely on monolithic rewriting by an LLM often degrade into shorter, less informative summaries over time, causing sharp performance declines (Figure 2). In domains such as interactive agents [trivedi2024appworld, patil2024gorilla, zhang2024caravan], domain-specific programming [ye2023generating, zhang2025adaptive], and financial or legal analysis [loukas2022finer, guha2023legalbench, wang2025finlora], strong performance depends on retaining detailed, task-specific knowledge rather than compressing it away.
@@
As applications such as agents and knowledge-intensive reasoning demand greater reliability, recent work has shifted toward saturating contexts with abundant, potentially useful information [jiang2025putting, chung2025long, chen2025flora], enabled by advances in long-context LLMs [peng2023yarn, mao2024lift]. We argue that contexts should function not as concise summaries, but as comprehensive, evolving playbooks—detailed, inclusive, and rich with domain insights. Unlike humans, who often benefit from concise generalization, LLMs are more effective when provided with long, detailed contexts and can distill relevance autonomously [jiang2025putting, liu2025selfelicit, suzgun2025dynamic]. Thus, instead of compressing away domain-specific heuristics and tactics, contexts should preserve them, allowing the model to decide what matters at inference time.
@@
To address these limitations, we introduce ACE (Agentic Context Engineering), a framework for comprehensive context adaptation in both offline settings (e.g., system prompt optimization) and online settings (e.g., test-time memory adaptation). Rather than compressing contexts into distilled summaries, ACE treats them as evolving playbooks that accumulate and organize strategies over time. Building on the agentic architecture of Dynamic Cheatsheet [suzgun2025dynamic], ACE incorporates a modular workflow of generation, reflection, and curation, while adding structured, incremental updates guided by a grow-and-refine principle. This design preserves detailed, domain-specific knowledge, prevents context collapse, and yields contexts that remain comprehensive and scalable throughout adaptation.
@@
We evaluate ACE on two categories of LLM applications that most benefit from comprehensive, evolving contexts: (1) agents [trivedi2024appworld], which require multi-turn reasoning, tool use, and environment interaction, where accumulated strategies can be reused across episodes; and (2) domain-specific benchmarks, which demand specialized tactics and knowledge, where we focus on financial analysis [loukas2022finer, wang2025finlora]. Our key findings are:
@@
    On the AppWorld benchmark leaderboard [AppWorldLeaderboard], ACE matches the top-ranked production-level agent IBM-CUGA [marreed2025towards] (powered by GPT-4.1) on average and surpasses it on the harder test-challenge split, while using a smaller open-source model (DeepSeek-V3.1).
@@ 2.1 Context Adaptation
Context adaptation (or context engineering) refers to methods that improve model behavior by constructing or modifying inputs to an LLM, rather than altering its weights. The current state of the art leverages natural language feedback [shinn2023reflexion, yuksekgonul2024textgrad, agrawal2025gepa]. In this paradigm, a language model inspects the current context along with signals such as execution traces, reasoning steps, or validation results, and generates natural language feedback on how the context should be revised. This feedback is then incorporated into the context, enabling iterative adaptation. Representative methods include Reflexion [shinn2023reflexion], which reflects on failures to improve agent planning; TextGrad [yuksekgonul2024textgrad], which optimizes prompts via gradient-like textual feedback; GEPA [agrawal2025gepa], which refines prompts iteratively based on execution traces and achieves strong performance, even surpassing reinforcement learning approaches in some settings; and Dynamic Cheatsheet [suzgun2025dynamic], which constructs an external memory that accumulates strategies and lessons from past successes and failures during inference. These natural language feedback methods represent a major advance, offering flexible and interpretable signals for improving LLM systems beyond weight updates.
@@ 2.2 Limitations of Existing Context Adaptation Methods
A recurring limitation of context adaptation methods is brevity bias: the tendency of optimization to collapse toward short, generic prompts. Gao et al. [gao2025prompt] document this effect in prompt optimization for test generation, where iterative methods repeatedly produced near-identical instructions (e.g., "Create unit tests to ensure methods behave as expected"), sacrificing diversity and omitting domain-specific detail. This convergence not only narrows the search space but also propagates recurring errors across iterations, since optimized prompts often inherit the same faults as their seeds. More broadly, such bias undermines performance in domains that demand detailed, context-rich guidance—such as multi-step agents, program synthesis, or knowledge-intensive reasoning—where success hinges on accumulating rather than compressing task-specific insights.
@@
In a case study on the AppWorld benchmark [trivedi2024appworld], we observe a phenomenon we call context collapse, which arises when an LLM is tasked with fully rewriting the accumulated context at each adaptation step. As the context grows large, the model tends to compress it into much shorter, less informative summaries, causing a dramatic loss of information. For instance, at step 60 the context contained 18,282 tokens and achieved an accuracy of 66.7, but at the very next step it collapsed to just 122 tokens, with accuracy dropping to 57.1—worse than the baseline accuracy of 63.7 without adaptation. While we highlight this through Dynamic Cheatsheet [suzgun2025dynamic], the issue is not specific to that method; rather, it reflects a fundamental risk of end-to-end context rewriting with LLMs, where accumulated knowledge can be abruptly erased instead of preserved.
@@ 3 Agentic Context Engineering (ACE)
We present ACE (Agentic Context Engineering), a framework for scalable and efficient context adaptation in both offline (e.g., system prompt optimization) and online (e.g., test-time memory adaptation) scenarios. Instead of condensing knowledge into terse summaries or static instructions, ACE treats contexts as evolving playbooks that continuously accumulate, refine, and organize strategies over time. Building on the agentic design of Dynamic Cheatsheet [suzgun2025dynamic], ACE introduces a structured division of labor across three roles (Figure 4): the Generator, which produces reasoning trajectories; the Reflector, which distills concrete insights from successes and errors; and the Curator, which integrates these insights into structured context updates. This mirrors how humans learn—experimenting, reflecting, and consolidating—while avoiding the bottleneck of overloading a single model with all responsibilities.
@@ 3.1 Incremental Delta Updates
A core design principle of ACE is to represent context as a collection of structured, itemized bullets, rather than a single monolithic prompt. The concept of a bullet is similar to the concept of a memory entry in LLM memory frameworks like Dynamic Cheatsheet [suzgun2025dynamic] and A-MEM [xu2025mem], but builds on top of that and consists of (1) metadata, including a unique identifier and counters tracking how often it was marked helpful or harmful; and (2) content, capturing a small unit such as a reusable strategy, domain concept, or common failure mode. When solving new problems, the Generator highlights which bullets were useful or misleading, providing feedback that guides the Reflector in proposing corrective updates.
@@ 4.1 Tasks and Datasets
    LLM Agent: AppWorld [trivedi2024appworld] is a suite of autonomous agent tasks involving API understanding, code generation, and environment interaction. It provides a realistic execution environment with common applications and APIs (e.g., email, file system) and tasks of two difficulty levels (normal and challenge). A public leaderboard [AppWorldLeaderboard] tracks performance, where, at the time of submission, the best system achieved only 60.3% average accuracy, highlighting the benchmark’s difficulty and realism.
@@
    Financial Analysis: FiNER [loukas2022finer] and Formula [wang2025finlora] test LLMs on financial reasoning tasks that rely on the eXtensible Business Reporting Language (XBRL). FiNER requires labeling tokens in XBRL financial documents with one of 139 fine-grained entity types, a key step for financial information extraction in regulated domains. Formula focuses on extracting values from structured XBRL filings and performing computations to answer financial queries, i.e., numerical reasoning.
@@ 4.2 Baselines and Methods
The base model is evaluated directly on each benchmark without any context engineering, using the default prompts provided by dataset authors. For AppWorld, we follow the official ReAct [yao2023react] implementation released by the benchmark authors, and build all other baselines and methods on top of this framework.
In-Context Learning (ICL) [agarwal2024many].
@@
MIPROv2 [opsahl2024optimizing].
MIPROv2 is a popular prompt optimizer for LLM applications that works by jointly optimizing system instructions and in-context demonstrations via bayesian optimization. We use the official DSPy implementation [DSPyMIPROv2], setting auto="heavy" to maximize optimization performance.
GEPA [agrawal2025gepa].
GEPA (Genetic-Pareto) is a sample-efficient prompt optimizer based on reflective prompt evolution. It collects execution traces (reasoning, tool calls, intermediate outputs) and applies natural-language reflection to diagnose errors, assign credit, and propose prompt updates. A genetic Pareto search maintains a frontier of high-performing prompts, mitigating local optima. Empirically, GEPA outperforms reinforcement learning methods such as GRPO and prompt optimizers like MIPROv2, achieving up to 10–20% higher accuracy with as much as 35× fewer rollouts. We use the official DSPy implementation [DSPyGEPA], setting auto="heavy" to maximize optimization performance.
Dynamic Cheatsheet (DC) [suzgun2025dynamic].
DC is a test-time learning approach that introduces an adaptive external memory of reusable strategies and code snippets. By continuously updating this memory with newly encountered inputs and outputs, DC enables models to accumulate knowledge and reuse it across tasks, often leading to substantial improvements over static prompting methods. A key advantage of DC is that it does not require ground-truth labels: the model can curate its own memory from its generations, making the method highly flexible and broadly applicable. We use the official implementation released by the authors [Suzgun2025_DynamicCheatsheet_code] and set it to use the cumulative mode (DC-CU).
@@
ACE optimizes LLM contexts for both offline and online adaptation through an agentic context engineering framework. To ensure fairness, we use the same LLM for the Generator, Reflector, and Curator (non-thinking mode of DeepSeek-V3.1 [deepseekai2024deepseekv3technicalreport]), preventing knowledge transfer from a stronger Reflector or Curator to a weaker Generator. This isolates the benefit of context construction itself. We adopt a batch size of 1 (constructing a delta context from each sample). We set the maximum number of Reflector refinement rounds and the maximum number of epochs in offline adaptation to 5.
@@ 4.3 Results on Agent Benchmark
Notably, on the latest AppWorld leaderboard (as of September 20, 2025; Figure 5), on average, ReAct + ACE (59.4%) matches the top-ranked IBM CUGA (60.3%), a production-level GPT-4.1–based agent [marreed2025towards], despite using the smaller open-source model DeepSeek-V3.1. With online adaptation, ReAct + ACE even surpasses IBM CUGA by 8.4% in TGC and 0.7% in SGC on the harder test-challenge split, underscoring the effectiveness of ACE in building comprehensive and self-evolving contexts for agents.
@@ 5 Discussion
Although ACE produces longer contexts than methods such as GEPA, this does not translate to linearly higher inference cost or GPU memory usage. Modern serving infrastructures are increasingly optimized for long-context workloads through techniques such as the reuse [gim2024prompt, yao2025cacheblend], compression [liu2024kivi, liu2024cachegen], and offload [lee2024infinigen] of KV cache. These mechanisms allow frequently reused context segments to be cached locally or remotely, avoiding repetitive and expensive prefill operations. Ongoing advances in ML systems suggest that the amortized cost of handling long contexts will continue to decrease, making context-rich approaches like ACE increasingly practical in deployment.
@@
Online and continuous learning are key research directions in machine learning for addressing issues like distribution shifts [koh2021wilds, gulrajani2021domain] and limited training data [pan2010survey, hutchinson2017overcoming, zhuang2019transfer]. ACE offers a flexible and efficient alternative to conventional model fine-tuning, as adapting contexts is generally cheaper than updating model weights [brown2020gpt3, lester2021prompttuning, li2021prefixtuning, hu2021lora]. Moreover, because contexts are human-interpretable, ACE enables selective unlearning [cao2015unlearning, bourtoule2021sisa, liu2024rethinkingllmunlearning]—whether due to privacy or legal constraints [gdpr2016art17, ccpa1798_105], or when outdated or incorrect information is identified by domain experts. These are promising directions for future work, where ACE could play a central role in advancing continuous and responsible learning.